Solve by substitution method.. Can you help me please?
2x+7y=6
x=10-7y

Answer by rfer(16322)

2(10-7y)+7y=6
20-14y+7y=6
20-7y=6
-7y=6-20
y=-14/-7
y=2
-------------
2x+7*2=6
2x+14=6
2x=-8
x=-8/2
x=-4
--------------
x=10-7y
-4=10-7*2
-4=10-14